OpenStudy (bananas1234):

Divide and simplify: (a-4)(a + 3)/a^2 - 2 divided by 2(a-4)/(a + 3)(a-1)

OpenStudy (bananas1234):

@jim_thompson5910

OpenStudy (bananas1234):

@robtobey

jimthompson5910 (jim_thompson5910):

When you divide two fractions, you will flip the second fraction and multiply Example: 1/2 divided by 3/4 = 1/2 times 4/3 So in this case, it means \[\Large \frac{(a-4)(a+3)}{a^2-2} \div \frac{2(a-4)}{(a+3)(a-1)}\] turns into \[\Large \frac{(a-4)(a+3)}{a^2-2} \times \frac{(a+3)(a-1)}{2(a-4)}\]

jimthompson5910 (jim_thompson5910):

I'll let you simplify

OpenStudy (bananas1234):

im not sure i know how to simplify can you walk me through it? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

well I can only see one pair of terms cancelling. Can you see it too?

OpenStudy (bananas1234):

the 2 (a + 3)

jimthompson5910 (jim_thompson5910):

no, the "(a-4)" \[\Large \frac{(a-4)(a+3)}{a^2-2} \times \frac{(a+3)(a-1)}{2(a-4)}\] \[\Large \frac{{\color{red}{(a-4)}}(a+3)}{a^2-2} \times \frac{(a+3)(a-1)}{2{\color{red}{(a-4)}}}\] \[\Large \frac{{\color{red}{\cancel{(a-4)}}}(a+3)}{a^2-2} \times \frac{(a+3)(a-1)}{2{\color{red}{\cancel{(a-4)}}}}\] \[\Large \frac{a+3}{a^2-2} \times \frac{(a+3)(a-1)}{2}\]

OpenStudy (bananas1234):

i see

jimthompson5910 (jim_thompson5910):

now multiply straight across

OpenStudy (bananas1234):

a+3^2/2a?

jimthompson5910 (jim_thompson5910):

more like \[\Large \frac{a+3}{a^2-2} \times \frac{(a+3)(a-1)}{2}\] \[\Large \frac{(a+3)(a+3)(a-1)}{2(a^2-2)}\] \[\Large \frac{(a+3)^2(a-1)}{2(a^2-2)}\]

OpenStudy (bananas1234):

a+3^2/2?

jimthompson5910 (jim_thompson5910):

you can't cancel anything else

OpenStudy (bananas1234):

so thats it? :)